OpenStudy (anonymous):

find the equation of the line that passes through the point A(1,-4,2) and is parallel to the intersection line of the two planes x-2y+3z-1=0 and x-4y+2z-8=0

OpenStudy (anonymous):

point 1,-4,2 ?!?! No can be.

OpenStudy (aroub):

gandal it can be!

OpenStudy (anonymous):

i got point (-4,-1/2,1) but i am not sure if i am right

OpenStudy (anonymous):

I can help you

OpenStudy (anonymous):

v₁ = {1,-2,3} ..... is a vector normal to x-2y+3z-1=0 v₂ = {1,-4,2} ..... is a vector normal to x-4y+2z-8=0 v₁ x v₂ = {8, 1, -2} ... is a vector parallel to the intersection line of the two planes ⇒ x = 1 + 8t ; y = -4 + t ; z = 2 - 2t .... is the desired line through (1,-4,2) parallel to {8, 1, -2} Answer: x = 1 + 8t ; y = -4 + t ; z = 2 - 2t

OpenStudy (anonymous):

Satisfied?

OpenStudy (anonymous):

WO

OpenStudy (anonymous):

thanks i just dont know why i;m getting this results